DIRC Operations Report
Wednesday 2008/03/05 to Wednesday 2008/03/12
Oncall: Jennifer Prendki (Jose Benitez on backup)
Fast Monitoring: Jennifer Prendki and Jose Benitez
Luminosities this week
PEP peak lumi: ~11.0 x 1033
(last week: ~9.5 x 1033)
PEP delivery: ~5.0 fb -1 (known bug)
BaBar integrated lumi: ~6.0 fb-1 (known bug)
Run 7 total delivery: 41.56 fb -1 (known bug)
Run 7 total BaBar recorded: 41.65 fb -1 (known bug) total,
2.89 fb-1 off-peak data
0.62 fb-1 on the Y(4S)
30.66 fb-1 on the Y(3S)
7.48 fb-1 on the Y(2S)
PEP activities/issues this week
- High backgrounds appeared through the week-end, mostly seen
by the DIRC and the LSTs.
BaBar activities/issues this week
- Y(2S) data taken this week.
DIRC activities/issues this week
- Recurrent scaler alarms starting through the week-end until
now (mostly sector 3). Threshold kept at it's current value to use
DIRC scaler alarm as a IFR background monitor.
- No other issue for the DIRC this week.
See DIRC logbook for details
Lumi and Background this week
- We are trying to correlate our scaler rates with the
background seen by the IFR people to see if they can be used to
monitor the LST backgrounds.
- Jennifer takes the pager next week. Jose is in back-up.